The instructions say to solve by factoring. x²+11x+10=0. In the notes we got two answers on similar questions. Please tell me how to solve!
Answer by ewatrrr(24785)

 
Hi,
x²+11x+10=0.
(x+10)(x+1) = 0
(x+10)= 0, x = -10
(x+1) = 0, x = -1